Mechanics

Two reversible engines are connected in series. The first one receives heat at T° K and rejects heat at 1527 'C. The second one receives the heat rejected by the first at 1527 'C and then expels heat at 527 'C. Find the temperature T for equal efficiencies of the two engines. (a) 3240' K (b) 2800' K (c) 2054' K (d) 1000' K (e) 3100' K

subhajit sarkar
7 Years agoGrade

1 Answer

APARNA SHAJI
7 Years ago
dear student am getting a value of 3250k                                                                                                                                              1800/T=800/1800 U VIL GET T= 4050 NOW SUBSTRACT 800 FROM IT AS IT IS IN SERIES CONNECTION. IF AM WRONG PLS CHECK YR OPTIONS AGAIN. HAPPY TO HELP U ALL THE BEST